ABC has employees 300 workers as a team. Standard production of 800 pieces per hour. Each employee in the team is paid a bonus for the excess production in addition to wages at hourly rates.

The bonus= % of production more than the standard quantity is found. Five % is regarded as the employees share. The employee in the team is paid as a bonus this percentage of a wage  rate of $30 per hours.

Total hours worked=     2100                            

Total Production=2700000

During the week Paul worked 40 hours and was paid $20 per hour. Margaret also worked 40 hours and was paid $20.

Calculate

     The bonus rate for the week                               

     Total bonus for the Team                                   

     Total pay for Paul and Margaret.                               

 

              

 



1
Expert's answer
2021-11-19T10:30:08-0500

Standard production for actual hours worked:

= Actual hours worked "\times" Standard production per hour

= 2,100 "\times" 1,000

= 2,100,000


Excess production:

= Actual production - Standard production

= 2,700,000 - 2,100,000

= 600,000


Computing percentage of excess production over standard:

"= \frac{600,000}{ 2,100,000}"


"= 0.2857"


Computing each employee's share of bonus:

= Percentage of excess production "\times 5\%"

= "28.57\%\times5\%"

= "0.1428"


Computing bonus rate:

= Hourly wage rate "\times" Men's share

="\$30 \times 0.1428"

= "\$4.284" per hour


Computing total amount of bonus for the week:

="2,100 \times \$4.284"

="\$8,996.4"

 

Computing total pay for Paul and Margaret:




Bonus rate for the week ="\$4.284" per hour


Total bonus for the team = "\$8,996.4" nearly "\$9,000"


Total pay for:

Paul = "\$971.36"

Margaret ="\$971.36"